Suppose a portfolio manager wishes to construct a portfolio using two securities Coll and USR, both of their return data are shown in the Slide 5 with the title “Hypothetical Investment Alternatives”. Specifically, the manger allocates $60,000 in the Coll and $30,000 in the USR. Please calculate (1) Portfolio Expected Return (2) Portfolio Standard Deviation

Economy

Prob.

T-Bills

HT

Coll

USR

MP

Recession

0.1

5.5%

-27.0%

27.0%

6.0%

-17.0%

Below avg

0.2

5.5%

-7.0%

13.0%

-14.0%

-3.0%

Average

0.4

5.5%

15.0%

0.0%

3.0%

10.0%

Above avg

0.2

5.5%

30.0%

-11.0%

41.0%

25.0%

Boom

0.1

5.5%

45.0%

-21.0%

26.0%

38.0%


Please do not use excel.
Answer #1

1) In order to calculate Portfolio Expected Return, we need to first calculate the expected return of the two securities i.e Coll and USR.

The expected return of Coll = 0.1*27+0.2*13+.4*0+0.2*(-11)+0.1*(-21)

= 2.7+2.6-2.2-2.1

=1%

The expected return of USR = 0.1*6+0.2*(-14)+0.4*3+0.2*41+0.1*26

= 0.6-2.8+1.2+8.2+2.6

=9.8%

The Expected Portfolio return = weight of Coll in portfolio* expected return of Coll + weight of USR in portfolio* expected return of USR

= 0.67*1+0.33*9.8

= 3.80%

2) In order to calculate portfolio standard deviation we need to calculate standard deviation of securities individually

Variance of Coll = 0.10(0.27-0.01)2 + (0.20)(0.13 - 0.01)2 + (0.40)(0.00 - 0.01)2   +(0.20)(-0.11 - 0.01)2+(0.10)(-0.21 - 0.01)2

   = 0.10*0.0676+0.20*0.0144+0.00004+0.20*0.0144+0.10*0.0484

=0.00676+0.00288+0.00004+0.00288+0.00484

= 0.0174

Variance of USR =0.10(0.06-0.098)2+0.20(0.14 - 0.098)2 + (0.40)(0.03 - 0.098)2   +(0.20)(0.41 - 0.098)2+(0.10)(0.26 - 0.098)2

        = 0.10*0.00144+0.0084+.00185+0.0194+0.0026

= 0.0324

Covariance between Coll and USR =0.10 (0.27-.01)(0.06-0.098)+0.20(0.13-0.01)(.14-0.098)+0.40(0.00-0.01)(0.03-0.098)+0.20(-0.11-0.01)*(0.41-0.098)+0.10(-0.21-.01)(0.26-0.098)

=-0.00098+0.00100+0.000272-0.00748-0.00356

=-0.010748

Portfolio Variance = w2A2(RA) + w2B2(RB) + 2*(wA)*(wB)*Cov(RA, RB)

=0.67*0.67*0.0174+0.33*0.33*0.0324+2*0.67*0.33*-0.010748

=0.00781+0.00353-0.00475

= 0.00659

The standard deviation is simply the square root of the variance therefore,

Portfolio Standard Deviation = 0.081 or 8.1%